OpenStudy (anonymous):

can someone check my answer :solve sqrt 2x-3=9 i got x=42

OpenStudy (saifoo.khan):

2x - 3 = 81

OpenStudy (anonymous):

yup

jimthompson5910 (jim_thompson5910):

Answer to \[\large \sqrt{2x-3}=9\] is x=42, good job

OpenStudy (anonymous):

OK - we now have an new expert on radicals !!! well done Jenni.

OpenStudy (anonymous):

squaring both sides x=42 ur answer is correct

OpenStudy (anonymous):

\[\sqrt{2x-3}=9\] \[\sqrt{2x-3}^2=9^2\] \[2x-3=81\] + 3 to both sides 2x=84 divide by 2 on both sides x=42

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

perfect.

OpenStudy (anonymous):

yay thanks for the help

OpenStudy (saifoo.khan):

Medals for Her.